show the coefficient for each reaction to balance the electrons lost = electrons gained
ITIN TIL Cileillilal Equation for the reaction that below. 2 MnO4-1 + 16 H+ + 5 C2042 -> 2 Mn+2 + 10 CO2 + 8 H20
Answer #1

anno 2002 mut2 + ca H-HA x+41-2)=-1| 2x+4(-2)=-2 X=42 X+21-2)=0 o to x=+7 2X= 76. x +4 23-13 oxidation half reaction Reductio